Status Report

ISS On-Orbit Status 10 May 2002

By SpaceRef Editor
May 10, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted
previously or below.

The crew continues to be well on top, or even ahead, of work
assignments, getting everything ready for the approaching UF-2
mission.

Early in the morning, CDR Yuri Onufrienko spent about 2.5 hours
attaching barcode labels to SM panel locations. [The labels,
required for use with the IMS BCR (inventory management system bare
code reader), were delivered by Progress-257 (7P). Yuri was to apply
the sticky-back labels to even and clean surface areas on the panel
interiors.]

The two flight engineers worked on the "shell game" of
preparing Lab and Node for returning two ZSRs (zero-G storage racks),
#187 & #115, on STS-111/UF-2 and receiving two new racks. [To
prepare for UF-2 deliveries, ZSR #112 was already location-swapped
with #115 yesterday.  Today’s work consisted of unloading
(continued from yesterday) and collapsing #187 in the Node and
replacing it with ZSR #110 (after installation of two K-BAR
adapters), then installing a protective shield behind it to protect
the bulkhead.  Later, Carl and Dan also unloaded ZSR #115,
transferring its storage items to ground-recommended locations.
Removal of #187 makes room for the up-coming EXPRESS Rack 3 (ER3),
while the arriving MSG (Microgravity Science Glovebox) replaces ZSR
#112 which in turn went into the position freed by #115.]

In further preparation for the arrival of ER3, Dan Bursch worked
on its future position (LAB1O3) to prevent structural interference
with its ARIS (active rack isolation system) operations.
 [The task included (1) removing upper and lower rack
closeout panels, (2) removing the standard rack electric grounding
strap with mounting bracket, and (3) modifying a standoff panel by
enlarging two notches in it with a ground-developed procedure using a
hand punch.]

On the ER2/ARIS rack Bursch also installed snubber alignment
guides to mechanically lock down the rack, in order to facilitate CDR
Yuri Onufrienko’s subsequent work on the ZCG (Zeolite crystal growth)
payload.

Onufrienko opened the ZCG furnace and removed all 19 autoclaves from
it.  Before that, he switched the primary and backup CPU
(central processor unit) circuit breakers off, as well as the power
to lockers 5 and 6. The activity was videotaped.

Onufrienko and Walz were advised to perform the planned swap of ZCG
hard drives during the autoclave unloading
activity because they both use the same tools. Also, with the furnace
cover removed, the swap-out was easier to accomplish.

FE-1 Walz also assisted the CDR in
another major (3-hr.) systems task: the disassembly and removal of
the KURS-A (active part of Russian rendezvous/docking radio) from the
Progress M1-8 (7P). Rather than discarding them with the cargo ship,
these rather expensive avionics elements are always returned to
Moscow for evaluation and recycling.

Dan Bursch continued his daily work on the BPS (biomass production
system) test facility (which earned him the title
"Agronaut" from POC).  [Today he again primed
(filled) the BPS NDS (nutrient delivery system) reservoir, transfer
pump and HCS (humidity control system) reservoir, while recording the
amount of air pulled from the tanks (if any). Later in the day, he
took gas samples from PGC (plant growth chamber) #2 and #4 and from
the cabin air, then swapped the root modules between PGC2 and PGC4.
During the chamber swapping, he was to comment on the condition of
the wheat plants (any yellowing, dead tips, elongated leaves, fungal
growth or any aromas), and to clean the interior of the chamber
covers from any water/debris.]

Today’s routine maintenance tasks were SOSH life support
systems inspection (Walz), 8A payload status check (Bursch), IMS
"delta" file preparation (Walz), IVCPDS (intravehicular
charged particle directional spectrometer) power-cycling for download
(Bursch), transfer of TVIS and RED exercise data to the MEC (Bursch),
and an IMS tag-up with Moscow via S-band (Onufrienko).

Late on 5/7, the crew inspected the TVIS to
study the SLD (subject load device) failure. When they found that the
right SLD wheel was not turning (the left SLD wheel turned OK) they
added some Braycote lubricant to the right wheel and were able to get
it to turn, although it is still noisy. MCC-H is assessing the crew’s
suggestion to possibly remove the wheel and re-lubricate it.

All crewmembers performed their daily 2.5-hr physical
exercise.

ISS navigational and motion control data continue to be supplied by
U.S. systems on the S0 truss: State vector (position/velocity) by
GPS-1, attitude by GPS-1 and RGA-1 (rate gyro assembly #1), and
attitudes rates by RGA-1.

CEO (crew earth observation) had the following targets today: E.
Mediterranean Dust and Smog
(as ISS approached the Sea of
Marmara and crossed the Dardanelles, the crew was to look left and
record any urban-industrial aerosols that may have been moving from
Istanbul westward toward the Aegean. From the Taurus Range on the
south coast of Turkey, they were to look right for plumes of Saharan
dust over Cyprus and the eastern Mediterranean. In addition to the
dust blow off Libya yesterday, another mass now appears to be
traveling NE-ward from Egypt)
, Oman – Tropical Cyclone
(Dynamic Event: The center of Tropical Cyclone 01A was just
left of the ground track as ISS exited the Arabian Peninsula. Crew
had an opportunity to photograph the internal structure of this
compact and imaginatively named storm. Good ESC [electronic still
camera] images that are downlinked will be candidates for posting to
Earth Observatory)
, Tropical Cyclone Kesiny (Dynamic
Event: Tropical Cyclone Kesiny had reached northern Madagascar by the
time of ISS overflight. The storm center was right of track, between
Madagascar and the Comoros Islands. Good ESC images that are
downlinked will be forwarded to the Earth Observatory website)
,
Congo-Zimbabwe Biomass Burning (conditions are clear over
south-central Africa, providing an opportunity to record any fires
and smoke. The pass began NW of Lake Mweru and continued across the
coast of Mozambique, where drought conditions have prevailed for
months)
, Angolan Biomass Burning (ISS approached SW
Africa near the Skeleton Coast and had a chance to record any smoke
or dust being borne out over the Atlantic by winds off the
continent)
, EasteNanited States (Dynamic Event:
Satellite images show fires burning in the Okeefenokee Swamp in the
SE corner of Georgia. Of interest: ESC images to the right of track,
toward the Swamp, then left of track to record NE-ward smoke
transport along the Atlantic coastal plain. Earth Observatory website
is interested in images of this event, as well)
, Sangre de
Cristo, NM Fires
(Dynamic Event: The pinon forests of the
Sangre de Cristo are tinderboxes following many dry months. As ISS
crossed the San Juan Mts. and approached the Sangre de Cristo, crew
was to look right and take ESC images of fires NE of Santa Fe)
,
Yucatan, C. America Smoke (Dynamic Event: Some reports this
morning suggest that fewer fires are now burning on the Yucatan
Peninsula; however, satellite images show a dense smoke blanket over
the region. Of interest: ESC images of the current situation, looking
right of track to document smoke movement over the Gulf of Campeche
and N toward Brownsville. As ISS traversed Honduras and Nicaragua,
crew was to try to record individual smoke sources along track, as
well as movement of smoke masses over the Caribbean and/or
Gulf)
.
CEO images can be viewed at the website http://eol.jsc.nasa.gov
<http://eol.jsc.nasa.gov/>

U.S. and Russian Segment Status (as of 1:18 pm
EDT)
:

Environmental Control and Life Support (ECLSS) and
Thermal Control (TCS):
Elektron O2 generator is Off (failed); SFOG candles ready
to be used.  Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5
(vacuum pump failed).  U.S. CDRA CO2 scrubber is Off.  BMP
Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in
Purify mode.

SM Working Compartment:  Pressure (mmHg) — 757,
temperature (deg C) — 26.0, ppO2 (mmHg) — 155.8, ppCO2 (mmHg) —
2.5 (?)
SM Transfer Compartment:  Pressure (mmHg) — 762,
temperature (deg C) — 19.4.
FGB Cabin:  Pressure (mmHg) — 756, temperature (deg C)
— 19.3.
Node:  Pressure (mmHg) — 749.61, temperature (deg C) —
21.9 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
U.S. Lab:  Pressure (mmHg) — 751.89, temperature (deg C)
— 24.0, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
Joint Airlock (Equip. Lock):  Pressure (mmHg) — 751.79,
temperature (deg C) — 22.4; shell heater temp (deg C) — 23.5, ppO2
(mmHg) — n/a; ppCO2 (mmHg) — n/a.
PMA-1:  Shell heater temp (deg C) — 22.2
PMA-2:  Shell heater temp (deg C) — 18.7

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not
available because MCA [major constituent analyzer] is failed and in
Extended Life mode
[= a state that preserves mass spectrometer
vacuum but produces no pp data]). MSA (mass spectrometer assembly)
and VGA (verification gas assembly) have been removed for return to
Earth.

Electrical Power Systems (EPS):
Beta Gimbal Assembly (BGA) 2B in Autotrack mode, BGA 4B in
Autotrack mode.
SM batteries: Battery #6 is offline (failed); battery #8 is in cycle
mode; all other batteries (6) are in “partial charge” mode.
FGB: Battery #1 is offline (storage mode); all other batteries (5)
are in “partial charge” mode.
Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby
mode.

Thermal Control Systems:
Air conditioner SKV-1 is On. SKV-2 is Off.

Command & Data Handling Systems:
C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1
is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is operational; PL-2 MDM on Standby.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Communications Systems:
All Russian communications & tracking systems are
nominal.
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
MCOR (medium-rate communications outage recorder) is operating
nominally.

Robotics:
SSRMS/Canadarm2 at Progress viewing position, with Keep Alive
power on both strings.
RWS (robotics workstations) are Off.
Troubleshooting continues on SSRMS Prime string Wrist Roll (WR)
joint.

ISS Orbit (as of this morning, 10:06 am EDT):
Mean altitude — 389.4 km
Apogee — 394.6 km
Perigee — 384.3 km
Period — 92.3 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0007581
Orbits per 24-hr. day — 15.59
Altitude decrease  315 m (mean) in last 24 hours
Solar Beta Angle:  -0.4 deg (magnitude decreasing)
Revolutions since FGB/Zarya launch (Nov. ’98) — 19816
Current Flight Attitude — LVLH (local vertical/local horizontal =
“earth-fixed”: z-axis in local vertical, x-axis in velocity vector
[yaw: -10 deg, pitch: -7.25 deg., roll: 0 deg]).

For more on ISS orbit and naked-eye visibility dates/times,
see

http://www.hq.nasa.gov/osf/station/viewing/issvis.html

SpaceRef staff editor.